# SegVol: Universal and Interactive Volumetric Medical Image Segmentation

### v3

@@ Abstract
Precise image segmentation provides clinical study with instructive information. Despite the remarkable progress achieved in medical image segmentation, there is still an absence of 3D foundation segmentation model that can segment a wide range of anatomical categories with easy user interaction. In this paper, we propose a 3D foundation segmentation model, named SegVol, supporting universal and interactive volumetric medical image segmentation. By scaling up training data to 90K unlabeled Computed Tomography (CT) volumes and 6K labeled CT volumes, this foundation model supports the segmentation of over 200 anatomical categories using semantic and spatial prompts. Extensive experiments on 10 internal validation tasks and 18 external validation tasks verify that SegVol outperforms the state of the art by a large margin. Through its capacity to provide precise volumetric segmentation across various anatomical categories, SegVol has the potential to accelerate advancements in medical imaging diagnosis and facilitate treatment optimization. The model and code are publicly available at: https://github.com/BAAI-DCAI/SegVol.
Figure 1: Overview of the model architecture and representative samples in the joint dataset. a. SegVol can model the 3D anatomical structures from volumetric inputs with easy user interactions including bounding box, point and text prompts. b. The joint dataset encompasses various anatomical structures in major regions of the human body. Several volume examples are demonstrated as 2D slices and 3D shapes in the images respectively.
@@ 1 Introduction
Volumetric image segmentation, involving extracting 3D regions of interest, such as organs, lesions, and tissues, plays a pivotal role in medical image analysis by accurately modeling the 3D structural information of the human body from volumetric medical images such as CT or MRI. This technique benefits numerous clinical applications including tumors monitoring[51, 62], surgical planning[41, 12], disease diagnosis[4], therapy optimization[52, 68], etc.
@@
Compared to 2D medical image segmentation[5, 72, 57, 67, 69, 42, 25, 70, 13], volumetric image segmentation is notably more challenging due to the labor-intensive annotation and resource-consuming computation. The research of volumetric medical image segmentation has garnered substantial attention, leading to a series of advancements[16, 71, 15, 61, 24, 33]. However, there exist several key limitations of the above-mentioned methods, which prevent their application in challenging tasks, e.g., liver tumor or colon cancer segmentation[2, 7, 43, 14], and real-world tasks, e.g., human-interactive segmentation[31, 38, 46, 23]. Firstly, the publicly available volumetric medical image datasets usually consist of a small number of mask annotations from a few varying categories. Due to the different label spaces, the traditional task-specific segmentation models trained on one dataset have difficulty in generalizing to others. For example, the CT-ORG dataset[47, 48, 2, 8] contains ‘lungs’ category, while this category is split into two sub-classes and named ‘left lung’ and ‘right lung’ in the LUNA16 dataset[54]. The main reason is that these models do not understand the semantics of anatomical categories. Secondly, traditional segmentation models have inferior performance when segmenting complex structures, such as tumors and cysts[27]. This is because these models are trained on insufficient data and are also not able to leverage the spatial information through user interaction. Thirdly, previous solutions are computationally expensive in the inference process. They typically employ a sliding window to infer the whole volumetric input. This strategy is not only time-consuming but also short-sighted, as the sliding window contains only local information. Recently, there have been some works[38, 6, 64] that introduce spatial prompts into medical image segmentation. However, most of them lack the ability to process the 3D input directly, e.g. [38, 6], and none of them is able to understand the semantics of anatomical categories.
Motivated by the success of 2D image analysis[31, 38], we present the first foundation model, SegVol, for volumetric medical image segmentation. SegVol enables universal and interactive segmentation of more than 200 anatomical categories, supporting both spatial and semantic prompts. SegVol is built on a lightweight architecture, ensuring its efficiency for practical medical image analysis. We summarize the key features of SegVol as follows:
@@
    Implement the massive pre-training on 96K CT volumes and the supervised fine-tuning on 25 public volumetric medical image segmentation datasets.
@@
    Enable semantic-prompt segmentation on over 200 anatomical categories by integrating the language model into the segmentation model.
@@
    Employ a synergistic mechanism to coordinate the spatial-prompt and semantic-prompt in the model and achieve high-precision segmentation.
@@
    Design a zoom-out-zoom-in strategy that significantly reduces the computational cost, meanwhile preserving precise segmentation.
We extensively evaluate the proposed SegVol on 10 internal validation tasks and 18 external validation tasks, which encompass a variety of anatomical structures including organs, tissues, and lesions. The internal validation experiments show that SegVol outperforms the traditional task-specific segmentation models, and the external validation experiments demonstrate that our method surpasses the state-of-the-art interactive models by a large margin. These experimental results verify the capabilities of SegVol as a foundation model for universal and interactive volumetric medical image segmentation.
Figure 2: Quantitative results and visualization of internal validation experiments. a. Violin plots for comparing experiment results of SegVol and task-specific methods. The vertical axis is the Dice score. b. The performance of SegVol improves as the training dataset scales up. c. The comparison of the average Dice score of SegVol and nnU-Net[24] across 3 internal lesion tasks. d. Visualization results of SegVol and nnU-Net across 3 internal lesion tasks.
2 Results
2.1 SegVol: a 3D foundation model for volumetric medical image segmentation
It is a long-standing challenge in medical image analysis to build a 3D foundation model that is capable of handling a wide range of segmentation tasks while achieving precise segmentation results. This challenge has two critical aspects. On the one hand, volumetric medical imaging datasets are usually dispersed and small in scale, making it challenging to train a comprehensive 3D foundation model that can generalize well across diverse datasets. On the other hand, segmentation tasks in medical image analysis encompass a wide range of semantic categories and spatial scales, ranging from organ to lesion segmentation, further complicating the development of a general solution.
To establish a universal volumetric segmentation model proficient in multiple tasks, we collect 25 CT volume segmentation datasets from public medical datasets and process them into a joint dataset, involving popular segmentation tasks. A total of 5,772 CT volumes of the joint dataset participate in the training and internal validation, with 149,199 volumetric masks and semantics. The collected joint dataset includes major regions of the human body, i.e., the head, neck, thorax, abdomen, and pelvis, comprising over 200 categories of organs and tissues, and 28 lesion tasks from different benchmarks. Some representative samples are shown in Fig. 1 b. The detailed information of joint dataset can be obtained from Supplementary Tables 1, 2, and Fig. 2.
Developing a universal segmentation model is challenging due to two main obstacles. One is the large number of categories leading to ambiguous semantics, where the same voxel may correspond to multiple targets. Another obstacle lies in the wide range of spatial scales of targets, varying from small lesions to large organs, and the complex and diverse structures of targets in the space. To address these challenges, our method adopts innovative strategies demonstrated in Fig. 1 a. As a universal model, our approach delivers accurate segmentation results for over 200 important organs, tissues, and lesions by leveraging text prompts to clarify semantic references. Furthermore, as a precise segmentation model, SegVol introduces point and bbox(bounding box) spatial prompts to guide the segmentation of anatomical structures, thus achieving high-precision segmentation performance. By leveraging these techniques, our method navigates the ambiguous semantics arising from numerous categories and accommodates the varying spatial scales and complex structures of targets, ensuring robust and precise segmentation across a wide range of medical imaging tasks.
2.2 Internal validation compared with task-specific segmentation models
Task-specific segmentation models mainly fall into two architectures, CNN-based models and Transformer-based models. We conduct internal comparative experiments with representative CNN-based models i.e. 3DUX-Net[33] and nnU-Net[24], and representative Transformer-based models i.e. SwinUNETR[15]. We conduct internal validation experiments on the test set of joint dataset, which is not observed during the training phase of SegVol. The 10 internal segmentation tasks are selected from BTCV[32] and MSD-spleen[58] datasets, which focus on organ segmentation, and from MSD-lung, MSD-colon, and MSD-liver datasets, which focus on lesion segmentation. We train task-specific segmentation models on each dataset individually for each method.
The quantitative experimental results are summarized in Fig. 2 a. Generally speaking, SegVol, jointly trained on 25 datasets, outperforms traditional task-specific segmentation models trained on a single dataset. Compared to these strong baselines, SegVol exhibits a narrower distribution of DSC scores across the eight tasks, indicating its robustness and good generalization ability. This mainly owes to the massive knowledge learned from diverse samples of the same categories but different datasets. SegVol depicts excellent performance on lesion tasks which are more challenging in semantic understanding and spatial locating. We present a detailed comparison to nnU-Net[24] on lesion tasks. As shown in Fig. 2 c, the average Dice score of SegVol is 14.76% higher than that of nnU-Net for lesion tasks. We visualize the prediction results of the two methods in Fig. 2 d, which intuitively show that SegVol performs more precise segmentation of the tumors than nnU-Net. The detailed scores and visualization results of interval validation are presented in Supplementary Table 3 and Fig. 3-5.
We analyze that there are mainly three factors that make SegVol more powerful than traditional task-specific models: 1) Massive generative pre-training on unlabeled data endows SegVol with a complete understanding of the volumetric structures and the discriminative feature representations, which is much superior to learning from a small number of samples. 2) Learning from joint datasets with semantic prompts makes SegVol generalize better to unseen data and categories. For instance, SegVol can learn from both ‘left kidney’ and ‘kidney’ categories based on their semantic correlation, while traditional task-specific models treat the two categories independently. 3) SegVol can be prompted with (spatial) points/bboxes, which provide a precise spatial reference, and (semantic) texts, which disambiguate the overlap of multiple categories in the same space. In contrast, traditional methods are not able to understand semantics. This ability enables SegVol to perform better than traditional methods in challenging tasks, e.g., segmenting lesions.

### v5

@@ Abstract
Precise image segmentation provides clinical study with instructive information. Despite the remarkable progress achieved in medical image segmentation, there is still an absence of a 3D foundation segmentation model that can segment a wide range of anatomical categories with easy user interaction. In this paper, we propose a 3D foundation segmentation model, named SegVol, supporting universal and interactive volumetric medical image segmentation. By scaling up training data to 90K unlabeled Computed Tomography (CT) volumes and 6K labeled CT volumes, this foundation model supports the segmentation of over 200 anatomical categories using semantic and spatial prompts. To facilitate efficient and precise inference on volumetric images, we design a zoom-out-zoom-in mechanism. Extensive experiments on 22 anatomical segmentation tasks verify that SegVol outperforms the competitors in 19 tasks, with improvements up to 37.24% compared to the runner-up methods. We demonstrate the effectiveness and importance of specific designs by ablation study. We expect this foundation model can promote the development of volumetric medical image analysis. The model and code are publicly available at: https://github.com/BAAI-DCAI/SegVol.
Figure 1: Overview of SegVol model architecture. SegVol produces precise segmentation of 3D anatomical structures from volumetric inputs with easy user interactions, including point, bounding box, and text prompts. Zoom-out-zoom-in mechanism: SegVol initially produces a rough prediction mask with zoom-out inference, then refines it with zoom-in inference on the identified ROI.
@@ 1 Introduction
Volumetric medical segmentation, involving extracting 3D regions of interest, such as organs, lesions, and tissues, plays a pivotal role in medical image analysis by accurately modeling the 3D structural information of the human body from volumetric medical images such as CT or MRI. The accurate segmentation can benefit numerous clinical applications including tumors monitoringsajid2019brain ; trebeschi2021prognostic , surgical planningminnema2022review ; ferrari2012value , disease diagnosischen2020deep , therapy optimizationsamarasinghe2021deep ; zaidi2010pet , etc.
Compared to 2D medical image segmentationchen2017deeplab ; zhou2018unet++ ; siddique2021u ; yin2022u ; zhang2018mdu ; mubashar2022r2u++ ; jha2020doubleu ; zhang2020dense ; gu2019net ; yang2024boostingweaklysupervisedreferringimage , volumetric image segmentation is notably more challenging due to the labor-intensive annotation and resource-consuming computation. Recently, the research of volumetric medical image segmentation has garnered substantial attention, leading to a series of advancementshatamizadeh2022unetr ; zhou2022nnformer ; hatamizadeh2022swin ; tang2022self ; isensee2021nnu ; lee20233d . However, existing volumetric medical segmentation methods have several key limitations which prevent their application in challenging tasks, e.g., liver tumor or colon cancer segmentationbilic2023liver ; christ2017automatic ; pacal2020comprehensive ; hamida2021deep , and real-world tasks, e.g., human-interactive segmentationkirillov2023segment ; ma2023segment ; ramadan2020survey ; hui2020linguistic ; bai2024m3dadvancing3dmedical .
@@
Firstly, the publicly available volumetric medical image datasets usually consist of a small number of mask annotations from a few varying categories. Due to the different label spaces, the traditional task-specific segmentation models trained on one dataset have difficulty in generalizing to others. For example, the CT-ORG datasetrister2019ct ; rister2018ct ; bilic2023liver ; clark2013cancer contains the ‘lungs’ category, while this category is split into two sub-classes and named ‘left lung’ and ‘right lung’ in the LUNA16 datasetsetio2017validation . Hence, a universal segmentation model has to understand the semantics of anatomical categories. Secondly, traditional segmentation models have inferior performance when segmenting complex structures, such as tumors and cystsjiang2022deep . This is because these models are trained on insufficient data and are also not able to leverage the spatial information through user interaction. Thirdly, previous solutions are computationally expensive in the inference process. They typically employ a sliding window to infer the whole volumetric input. This strategy is not only time-consuming but also short-sighted, as the sliding window contains only local information. Recently, there have been some worksma2023segment ; cheng2023sam ; wang2023sammed3d that introduce spatial-prompt into medical image segmentation, shown in Table 1. However, most of them lack the ability to process the 3D input directly and naturally, and none of them is able to understand the semantics of anatomical categories.
In this paper, we propose the first foundation model for volumetric medical image segmentation – SegVol. The proposed model enables universal and interactive 3D segmentation of more than 200 anatomical categories, supporting both spatial and semantic prompts. SegVol can also be driven by the combination of multi-prompt, like ‘bounding box+text’ or ‘point+text’ prompts, achieving high-precision segmentation and semantic disambiguation. To enable efficient and precise segmentation of volumetric images, we develop a zoom-out-zoom-in mechanism that enables the model to be efficient and precise. We evaluate the proposed SegVol on 22 volumetric medical image segmentation tasks and the results demonstrate our method surpasses other SAM-like interactive segmentation methodskirillov2023segment ; cheng2023sam ; wang2023sammed3d ; ma2023segment by a large margin. Extensive case studies and ablation experiments are also carried out to prove the advantages of SegVol and the effectiveness of the zoom-out-zoom-in mechanism and multi-prompt combination.
We summarize our key contributions as follows:
@@
    Collect and process 25 public volumetric medical segmentation datasets, encompassing over 200 anatomical categories. The pseudo label is introduced to relieve the spurious correlation in the training data.
@@
    Implement massive 3D pre-training on 96K CT volumes and supervised fine-tuning on the 6k labeled datasets.
@@
    Support spatial-prompt, semantic-prompt, and combined-prompt segmentation, achieving high-precision segmentation and semantic disambiguation.
@@
    Design a zoom-out-zoom-in mechanism that significantly reduces the computational cost, meanwhile preserving precise segmentation.
2 Methodology
2.1 Dataset Construction
One of the main challenges of training a universal volumetric medical segmentation model is the absence of large-scale publicly available volumetric medical data, especially CTs with segmentation annotations. Doing our utmost, we collected 25 open-source segmentation CT datasets, including CHAOSCHAOS2021 ; CHAOSdata2019 ; kavur2019 , HaN-Segpodobnik2023han , AMOS22ji2022amos , AbdomenCT-1kMa-2021-AbdomenCT-1K , KiTS23heller2023kits21 , KiPA22he2021meta ; he2020dense ; shao2011laparoscopic ; shao2012precise , KiTS19heller2020state , BTCVlandman2015miccai , Pancreas-CTroth2016data ; roth2015deeporgan ; clark2013cancer , 3D-IRCADBsoler20103d , FLARE22simpson2019large ; ma2021abdomenct , TotalSegmentatorwasserthal2022totalsegmentator , CT-ORGrister2019ct ; rister2018ct ; bilic2023liver ; clark2013cancer , VerSe19, VerSe20sekuboyina2021verse ; loffler2020vertebral ; liebl2021computed , SLIVER07heimann2009comparison , QUBIQQUBIQ_2021 , six MSD datasetssimpson2019large , LUNA16setio2017validation , and WORDluo2022word . Their detailed information and availability are shown in the Section A. These CTs originate from various medical institutions, captured by different machines with varying parameter settings and scanning regions. To standardize these datasets, we use the mean voxel value of each volume to filter the background and then perform normalization on the foreground voxels.
Volumetric segmentation datasets suffer from the notorious problem of partial labels. Most of these datasets have annotations of only a few segmentation targets, e.g., several organs. Therefore, the deep models may learn the spurious correlation between datasets and segmentation targets, and thus produce inferior results during the inference phase. To relieve this problem, we introduce the pseudo labels by utilizing the Felzenswalb-Huttenlocher (FH)felzenszwalb2004efficient algorithm to generate pseudo masks for each CT scan. Pseudo masks can supplement unlabeled categories in a dataset, therefore relieving the spurious correlation problem. To restrain the noise and numerous tiny masks in pseudo labels, we employ the following strategies: 1) The pseudo masks are replaced with ground truth masks when applicable. 2) We filter out tiny structures smaller than 1‰ of the whole volume size. 3) Each mask is refined by dilation and erosion operations.
Table 1: The different settings and functions of SAM-like interactive segmentation methods.
Prompt Type
Method Image Domain Dimension Training Point Bbox Text Inference Input
SAMkirillov2023segment Natural 2D Full-Param 1024$\times$1024
MedSAMma2023segment Medical 2D Decoder 1024$\times$1024
SAM-Med2Dcheng2023sam Medical 2D Adapter 1024$\times$1024
SAM-Med3Dwang2023sammed3d Medical 3D Full-Param 128$\times$128$\times$128
OURS Medical 3D Full-Param Full Resolution
2.2 Model Architecture
Motivated by the recent advance in 2D nature image segmentation, Segment Anything (SAM)kirillov2023segment , we design a novel model for interactive and universal volumetric medical image segmentation, named, SegVol. The model is illustrated in Figure 1. SegVol supports three types of prompts for interactive segmentation: ‘bounding box(bbox)’ prompt, including the coordinates of two diagonal vertices; ‘point’ prompt, composed of a set of positive and negative points; and ‘text’ prompt, such as ‘liver’ or ‘cervical spine C2’. The model consists of four modules: image encoder, text encoder, prompt encoder, and mask decoder.
We employ 3D ViT (Vision Transformer)dosovitskiy2021image ; hatamizadeh2021unetr as the image encoder, which exhibits remarkable advantages over convolutional modelshe2016deep when pre-trained on large-scale datasets. The 3D ViT structure is designed as follows: patch size=(4, 16, 16), layers number=12, heads number=12, hidden size=768. We first pre-train 3D ViT using SimMIM algorithmxie2022simmim on the collected 96K CTs, and then conduct further supervised fine-tuning on the 6K CTs with 150K labeled segmentation masks.
One of the main limitations of traditional segmentation models is that the models learn dataset-specific labels encoded as integers which cannot generalized to unseen datasets or tasks, preventing their real-world applications. We enable universal segmentation across datasets by leveraging the text encoder from CLIP modelradford2021learning to encode the input text prompt, as CLIPradford2021learning has been trained to align image and text embeddings on web-scale image-text pairs. Given a word or phrase as the text prompt, we complete it using the template ‘A computerized tomography of a [text prompt]liu2023clipdriven and then encode it into text embedding. The off-the-shelf text encoder is frozen during training due to the limited text data in CT datasets. Following SAMkirillov2023segment , we obtain the spatial-prompt embedding using positional encodingtancik2020fourier on point and bbox prompt.
After obtaining the image embedding and prompt embedding, we input them into the mask decoder and predict the mask. We use self-attention and cross-attention in two directions to fuse the image embedding and prompt embedding, and then employ the transposed convolutions and interpolation operations to generate masks. Since text embedding is the key to universal segmentation and it is also challenging to learn the correlation between text and volumetric regions, we enhance the text information by introducing a parallel text input branch beside the joint prompt embedding.
2.3 Prompt Generation
SegVol accepts multiple types of prompts, including individual point, bbox, and text prompts, and also their combinations. To make full use of the segmentation training data, we generate kinds of prompts for each datum and construct kinds of prompt-mask data pairs for training.
The point prompt is built from ground truth or pseudo masks, consisting of three kinds of points, namely, positive point, negative point, and ignored point. Positive point means that it is within the target mask region, while negative points are those outside. Ignored points are utilized to ensure a uniform length of the point prompts for input completion. Notably, these ignored points are not considered by the model.
The text prompts are constructed based on their category names. As pseudo masks produced by the unsupervised FH algorithmfelzenszwalb2004efficient do not have the semantic information, we only use point and bbox prompts for training on masks of pseudo labels.
2.4 Zoom-out-zoom-in Mechanism
SAM-like interaction with large-volume images is laborious for users, especially in the scene where the sliding window has to be used due to the limited view of those 3D models. To provide users with an easy SAM-like interface, we design a zoom-out-zoom-in mechanism, which is efficient and precise, consisting of zoom-out-zoom-in inference and multi-size training. As demonstrated in Figure 1, the zoom-out process involves resizing a volumetric image, which is input into the model with user prompts to generate a coarse segmentation mask. Then, the Region of Interest (ROI) from the original image is cropped for zoom-in analysis. In the zoom-in process, a sliding window is used to perform precise inference driven by prompts generated from the coarse segmentation mask. After that, the ROI prediction mask will be back-filled to the coarse segmentation mask to finish the final prediction. Besides, multi-size training involves augmenting the input data by resizing CTs for the zoom-out view and cropping them into cubes for the zoom-in view. The zoom-out-zoom-in mechanism realizes the computational cost reduction meanwhile producing precise segmentation of the ROI.
2.5 Loss Function
We apply SimMIM algorithmxie2022simmim to pre-train the image encoder of SegVol with the masked image modeling loss $\mathcal{L}_{\text{pre-training}}(\bm{\theta}_{\text{IE}};\mathcal{D}_{1})$. The loss function is as follows:
$\mathcal{L}_{\text{pre-training}}(\bm{\theta}_{\text{IE}};\mathcal{D}_{1})=\frac{1}{\Omega(\bm{a}_{\text{M}})}||\bm{b}_{\text{M}}-\bm{a}_{\text{M}}||_{1},$ (1)
where $\bm{\theta}_{\text{IE}}$ is the parameter set of SegVol’s image encoder. $\bm{a},\bm{b}\in\mathbb{R}^{D\times H\times W}$ are the input voxel values and predicted values, respectively. M denotes the set of masked voxels, $\Omega(\cdot)$ is the number of elements, and $\mathcal{D}_{1}$ is the pre-training dataset.
We combine the Binary Cross-Entropy (BCE) loss and Dice loss as the supervised fine-tuning loss function $\mathcal{L}_{\text{fine-tuning}}(\bm{\theta};\mathcal{D}_{2})$ to train the model with trainable parameters $\bm{\theta}$ (text encoder frozen). $\mathcal{D}_{2}$ is the supervised fine-tuning dataset and $\bm{x},\bm{y}\in\mathbb{R}^{D\times H\times W}$ are the predicted mask and ground-truth mask, respectively. $\mathcal{F}(\cdot,\bm{\theta})$ is the forward function of SegVol. The loss function is as follows:
$\mathcal{L}_{\text{BCE}}(\bm{\theta};\mathcal{D}_{2})=-\mathbb{E}_{(\bm{x},\bm{y})\sim\mathcal{D}_{2}}[\langle\bm{y},\log(\mathcal{F}(\bm{x},\bm{\theta}))\rangle+\langle 1-\bm{y},\log(1-\mathcal{F}(\bm{x},\bm{\theta}))\rangle]$ (2)
$\mathcal{L}_{\text{Dice}}(\bm{\theta};\mathcal{D}_{2})=1-\mathbb{E}_{(\bm{x},\bm{y})\sim\mathcal{D}_{2}}[\frac{2\cdot\langle\bm{y},\mathcal{F}(\bm{x},\bm{\theta})\rangle}{\|\bm{y}\|_{1}+\|\mathcal{F}(\bm{x},\bm{\theta})\|_{1}}]$ (3)
The detailed fine-tuning algorithm of SegVol is presented in Section B.
3 Experiments
In this section, we conduct extensive experiments on 22 volumetric medical image segmentation tasks to compare SegVol with other SAM-like medical image segmentation methodskirillov2023segment ; cheng2023sam ; wang2023sammed3d ; ma2023segment . Ablation studies are also carried out to prove the effectiveness of the zoom-out-zoom-in mechanism and provide more insights about dataset scale and multi-prompt combination. Detailed case studies are conducted to discuss the disambiguation ability of semantic-prompt and the capability of identifying the segmentation results with spatial-prompt.
3.1 Experimental Setup
During the pre-training, we follow SimMIM algorithmxie2022simmim to train the 3D ViT encoder of SegVol on the collected 96K CTs for 2000 epochs. In the supervised fine-tuning stage, we train SegVol (with the text encoder frozen) on the labeled 25 volumetric medical image segmentation datasets for 270 epochs with batch size 32 and input size (32, 256, 256), using AdamW optimizerloshchilov2019decoupled . SimMIM pre-training takes about $20\times 8$ GPU hours, while fine-tuning takes about $300\times 8$ GPU hours. All the above training process is implemented on 8 NVIDIA A100-SXM4-40GB. Three external datasetsji2022amos ; uls23 ; lambert2019segthor and 20% testing data preserved from 25 collected datasets are used in the following experiments.
Table 2: Quantitative comparative experiment results for SegVol and other 5 SAM-like interactive segmentation methods settings in terms of the median value of Dice score.
Dataset Category SAM(Point) kirillov2023segment SAM(Bbox) kirillov2023segment SAM-MED2D cheng2023sam SAM-MED3D wang2023sammed3d MedSAM ma2023segment OURS
AMOS22 ji2022amos Aorta 0.7267 0.4362 0.8704 0.8102 0.3387 0.9273
Bladder 0.4162 0.6281 0.8417 0.4338 0.6799 0.9120
Duodenum 0.1554 0.3192 0.5066 0.3820 0.3066 0.7402
Esophagus 0.2917 0.3541 0.5500 0.5174 0.3610 0.7460
Gallbladder 0.2831 0.6161 0.7999 0.5643 0.6609 0.8763
Adrenal gland(L) 0.0555 0.4222 0.5068 0.4584 0.3766 0.7295
Left kidney 0.8405 0.8274 0.9325 0.8723 0.7909 0.9489
Liver 0.7477 0.5124 0.6904 0.8801 0.6137 0.9641
Pancreas 0.2127 0.3392 0.5656 0.5391 0.3217 0.8295
Postcava 0.2042 0.5251 0.4436 0.6683 0.5211 0.8384
Prostate uterus 0.2344 0.6986 0.7518 0.6231 0.7739 0.8557
Adrenal gland(R) 0.0452 0.3642 0.1681 0.3708 0.3855 0.6994
Right kidney 0.8459 0.8215 0.9077 0.8632 0.7851 0.9505
Spleen 0.5936 0.6536 0.9267 0.8591 0.7038 0.9589
Stomach 0.4229 0.3883 0.5399 0.4576 0.4378 0.9123
Average 0.4050 0.5271 0.6668 0.6200 0.5371 0.8593
ULS23 uls23 DeepLesion3D 0.3686 0.7473 0.3258 0.2386 0.7680 0.7065
BoneLesion 0.4461 0.6671 0.1947 0.4447 0.6896 0.6920
PancreasLesion 0.0675 0.5579 0.5548 0.5526 0.6561 0.7265
Average 0.2941 0.6574 0.3584 0.4120 0.7046 0.7046
SegTHOR lambert2019segthor Aorta 0.2744 0.3894 0.8077 0.7703 0.3278 0.8439
Esophagus 0.0348 0.2046 0.3578 0.6394 0.2196 0.7201
Heart 0.6695 0.8876 0.6012 0.8325 0.8924 0.8172
Trachea 0.9147 0.1611 0.8306 0.8485 0.1261 0.8807
Average 0.4734 0.4107 0.6493 0.7727 0.3915 0.8155